one month Raina rented 2 movies and 3 video games for a total of $24. The next month she rented 6 movies and 5 video games for a total of $50. find the total cost for each movie and video game.
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39618)

simple two-variable linear system
very common exercise

Answer by ikleyn(52797)

.
one month Raina rented 2 movies and 3 video games for a total of $24.
The next month she rented 6 movies and 5 video games for a total of $50.
find the total cost for each movie and video game.
~~~~~~~~~~~~~~

Write equations as you read the problem


    2M + 3V = 24      (1)

    6M + 5V = 50      (2)


Multiply first equation by 3 (both sides).  Keep the second equation as is


    6M + 9V = 72      (1')

    6M + 5V = 50      (2')


From equation (1'), subtract equation(2')


          4V = 72 - 50 = 22

           V           = 22/4 = 5.50.


Then from equation (1)

    2M + 3*5.50 = 24

    2M          = 24 - 17.50 = 6.50

     M                       = 6.50/2 = 3.25.


ANSWER.  $3.25 for movie;  $5.50 for video-game.

Solved.

-------------

To see a million other similar  (and different)  solved problems,  look into the lessons
    - Roses and vilolets
    - Counting calories and grams of fat in combined food
    - Typical word problems on systems of 2 equations in 2 unknowns
    - OVERVIEW of lessons on solving systems of two linear equations in two unknowns

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Systems of two linear equations in two unknowns".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.